Plurals and Anaphora

Jeroen Groenendijk

(editor)

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852

Deliverable R2.2.A, Part I August 1993

Plurals and Anaphora

Jeroen Groenendijk Universiteit van Amsterdam

(editor)

authors:

Jaap van der Does Dorit Abusch

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852

Deliverable R2.2.A, Part I August 1993 esprit Basic Research Action 6852 dyana-2

Dynamic Interpretation of Natural Language

Universiteit van Amsterdam Institute for Logic, Language and Computation (ILLC)

University of Edinburgh Centre for Cognitive Science (ECCS)

Universität München Centrum für Informations- und Sprachforschung (CIS)

Universitetet i Oslo Department of Linguistics and Philosophy (ILF)

Universität Stuttgart Institut für Maschinelle Sprachverarbeitung (IMS)

Universität Tübingen Seminar für Sprachwissenschaft (SfS)

Rijksuniversiteit Utrecht Research Institute for Language and Speech (OTS)

For copies of reports, updates on project activities and other dyana-related information, contact:

The dyana-2 Project Administrator ILLC/Department of Philosophy University of Amsterdam Nieuwe Doelenstraat 15 NL-1012 CP Amsterdam

\begin{matrix}c \\ ◯ \end{matrix} 1993, The Individual Authors

No part of this document may be reproduced or transmitted in any form, or by any means, electronic or mechanical, including photocopy, recording, or any information storage and retrieval system, without permission from the copyright owner.

Contents

Task 2.2, subtask 1: Generalized Quantifiers, Plurals and Discourse Anaphora The Dynamics of Sophisticated Laziness Jaap van der Does . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Task 2.2, subtask 2: The scope of Indefinites Dorit Abusch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

Task 2.2, subtask 1

Generalized Quantifiers, Plurals and Discourse Anaphora

The Dynamics of Sophisticated Laziness

Jaap van der Does

(University of Amsterdam)

Task 2.2, subtask 2 The scope of Indefinites

The scope of Indefinites

Dorit Abusch

(Stuttgart University)